Find the cube root of each of the following numbers by prime factorisation method.

 

(ix) 175616

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(ix) 175616

Solution:

175616 = 2×2×2×2×2×2×2×2×2×7×7×7

By grouping the factors in triplets of equal factors,

175616 = (2×2×2)×(2×2×2)×(2×2×2)×(7×7×7)

Here, 175616 can be grouped into triplets of equal factors,

∴ 175616 = (2×2×2×7) = 56

Hence, 56 is cube root of 175616.


Thanks